OpenStudy (anonymous):

Can someone please help? In a company, the amount of "V"acation (in hours) an employee earns varies directly with the amount of "T"ime (in weeks) worked. An employee who works 2 weeks can earn 3 hours of vacation. a.) Write a direct variation equation that relates t and v. b.) How many hours of vacation time does an employee earn in 8 weeks?

OpenStudy (nikato):

if V varies directly with T then for a. it is v=kt where k is a constant/ then for b, using the equation in a, v=kt. you want to solve for constant k by plugging the given 2 for t and 3 for v so 3=2k divide both sides by 2 k=3/2 so your eqation for this problem is v=(3/2)t and for b, you want find many hours earned in 8 weeks so t=8 plug it into equation v=(3/2)8 and now multiply and solve for v